OpenStudy (anonymous):

Which of the following verbal expressions describes the algebraic expression x+6/x-8 ? A. eight less than a number divided by six more than a number B. the quotient of six more than a number and eight less than a number. C. the quotient of a number and a number, and the quotient of six and eight D. the quotient of six more than a number and the difference of eight and the number

OpenStudy (anonymous):

\[ \text{A describes something like...}\\ \ \ \ (n-8)\div(n+6)\\ \text{Conversely, B describes...}\\ \ \ \ \frac{n+6}{n-8}\\ \text{C on the other hand...}\\ \ \ \ \ \frac{n}n+\frac68\\ \text{And ultimately D corresponds to...}\\ \ \ \ \ \frac{n+6}{|n-8|} \]